calculate the morecular formula of a hydracarbon which contains 85.7percent of carbo and has molecular mass 84 grams

Grade:9

1 Answers

Vikas TU
7 years ago
Let the hydrocarbon be CxHy.
then given,
12x + y = 84....................(1)
and
12x/84 x 100 = 85.7
on solving,
x = 5.99 = 6 (approx.)
Put x in eqn. (1)
we get,
y = 12
Hence the Hydrocarbon is:
C6H12.